OpenStudy (anonymous):

Two consecutive positive integers have the property that one integer times twice the other equals 612. What is the sum of these two integers?

OpenStudy (anonymous):

i guess but i stilll no understand

OpenStudy (anonymous):

Ok one integer is represented by x. The consecutive integer would be x+1. do you understand that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so if x is the first and x+1 is the second, then 2x+2 is twice the second right?

OpenStudy (anonymous):

right and did you factor out a 2?

OpenStudy (anonymous):

well not yet...

OpenStudy (anonymous):

So the question reads that the first integer multiplied by 2 times the second will be 612. So x * (2x+2) = 612 Follow so far?

OpenStudy (anonymous):

im with you on this

OpenStudy (anonymous):

Great! So when you multiply out, you get 2x^2 + 2x = 612

OpenStudy (anonymous):

Then you divide everything by 2: x^2 + x = 306

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then you solve for x by getting it to the form ax^2 + bx + c =0 and factoring

OpenStudy (anonymous):

(x +18)(x - 17) = 0 x = -18, 17 But you need the two that are consecutive, and since there was a square, you can look for different signed answers. So the two integers are 17 and 18. To check: 17 * (2*18) = 17 * 36 = 612

OpenStudy (anonymous):

oh ok thanks now i understand it.

OpenStudy (anonymous):

great!